OpenStudy (anonymous):

A water-tank has a base with dimensions 2 feet by 6 feet. If a cube with each side 1foot is totally immersed in the water, how many inches will the water rise? (12inches= 1 foot)

OpenStudy (shane_b):

Let's use inches. The area of the base is 1728 in^2. You're adding 1728in^3...so it will rise 1 inch.

OpenStudy (shane_b):

I assume you know how I got those values for volume and area but if not just ask.

OpenStudy (anonymous):

no i don't remember the formulas

OpenStudy (anonymous):

also what is meant by 2 feet by 6 feet?

OpenStudy (shane_b):

\[Area_{rect}=L*W=(24")(72")=1729in^2\]\[Volume_{cube}=s^3=12"^3=1728in^3\]Feet is a unit of length which contains 12 inches per foot. It's giving you the measurements in feet which you should convert to inches before doing the calculations. I did that part for you.

OpenStudy (shane_b):

err 1728in^2 should have been that first number...I mistyped it.

OpenStudy (shane_b):

To be clear: 2ft = 24inches and 6 foot = 72 inches.

OpenStudy (anonymous):

why you have taken cube of 12

OpenStudy (shane_b):

Because that's how you obtain the volume of a cube...side * side * side = Volume.

OpenStudy (anonymous):

but 12 is not the length of 1 side

OpenStudy (shane_b):

In the question, each side of the cube is 1ft. 1ft = 12 inches. I converted everything to inches because it makes the problem simpler.

OpenStudy (anonymous):

sorry i didn't read the question carefully

OpenStudy (shane_b):

No problem...do you understand everything else now?

OpenStudy (anonymous):

yes